You are here: Foswiki>PartnerObserving Web>SDTTeamPage>InstrumentIssues (21 Jul 2009, NormCushing)Edit Attach

Issues that LBC Observers Should Know About

Last update: 7 February 2008

Data Analysis

1. Prior to the February 2007 SDT time, the exposure times in the headers had small offsets due to some telemetry problems with the shutter. The data from January 2007 SDT has had a script run on them to correct the headers.
In general, the commanded exposure times are the actual exposure times. Data taken in February through June 2007 should be fine.
Blue Shutter was rebuilt in August 2007, and Red Shutter was installed in November 2007. The final timing calibration for both shutters will be checked in January.

OB Preparation

1. From Olga Kuhn: You need to make OB's for focus/collimation if your field doesn't have suitable stars. Here's her email:

As the February SDT is starting, I wanted to raise again the issue of focus/collimation before observing extended source targets.

The usual method to focus and collimate the mirror is to slew to the target position and obtain an extra-focal image which is then fed to an IDL routine that automatically detects and analyzes the pupil images. This is compromised when it selects small but resolved sources like HII regions or distant galaxies. It can be time-consuming for the observer to find regions reasonably close to the target which are free of these resolved sources. See the discussion under "NOTE for Extended Source Targets" on the wiki for more information: http://wiki.lbto.arizona.edu/twiki/bin/view/PartnerObserving/OBPrepConsiderations

It would be most efficient if a region near the target, with stars but without many resolved sources, were pre-selected and the coordinates of it entered into the OB used to obtain the extrafocal image. I am attaching a template of this OB, "Bfastextra.ob" and wondered whether investigators who have extended source targets would either edit in, under TEC_RA and TEC_DEC, the coordinates (decimal degrees) of such a region (replace the values 0 and -90 in lines 14 and 17) or send us these coordinates ( observe@lbto.arizona.edu): one set per field. Please rename the "Bfastextra.ob" with a name which associates it easily to the science OB (e.g. Bfastextra_<obname>.ob)

The extra-focal pupil image is centered at the coordinates TEC_RA and TEC_DEC and uses only a 2048 pxl (7.85') x 1048 pxl (5.4') subregion of chip 2.

There are some tools such as ESO/SkyCat and GAIA which can overlay catalogs (USNO, GSC) on DSS images and can help with selection of a region.

PartnerObserving

1. Use V-BESSEL extrafocal for all Blue focussing, and use r-SLOAN extrafocal for all Red focussing. Do not focus with extrafocal images from other filters, because they have uncalibrated spherochromatism.

2. There are still some soft hangs of the LBC UI that often cure themselves (and rarely last more than 6 seconds). Multiple LBC Control windows are now permitted, but take care to issue commands from only one at a time. Try changing UI pages on your browser to free things up. (It appears that this is more of a Firefox problem than an lbc UI problem.)

3. LBC has a new feature so that if the "lbc" server program crashes, it automatically restarts and returns to the previous power state of the camera. This feature is your friend and makes restarting the camera significantly easier and faster, but it can also bite you if the program crashes in an unusual power state (it tries to return to that unusual state). (Delete the power.conf file if you need to cancel the remembered state.)

4. Normal protocol would be turn turn off the LBC camera(s) at the end of the night, and turn them on again in late afternoon. Leave housekeeping power on all the time. Be sure to turn the cameras off (including housekeeping) if thunderstorms threaten. However, it seems that the system is more stable and reliable if the cameras are left on 24 hours a day - this is acceptable practice in good weather. The serial protocol of the housekeeping telemetry seems to get confused if it is on continuously for several days. So if the camera hasn't been turned off for some other reason, you may need to cycle housekeeping off/on to restore the temperature and vacuum values.

5. The observer workstations, lbtmu03, lbtmu05, have a propensity to become very slow so that window and cursor management become impossible. The apparent cause is heavy memory use by Firefox causing swapping. Be careful about having Firefox using more than 20% of physical memory. See Issue #806 for details. We've set the sdtobs Firefox profiles to Clear Private Data (under Preferences - Privacy) each time Firefox is closed. If you close all your Firefox windows and restart Firefox once a day, everything should be happy. Note that killing Firefox can be done at anytime without interrupting observations. Note also that you must kill all of your Firefox windows, not just the one with the LBC UI.

6. Filter focus offsets for Blue are fairly well measured. Accurate filter focus offsets for Red were measured once during Red Commissioning time in January 2008. We could use some additional measurements of the Red filter focus offsets in good seeing. Beware that the apparent focus offset can change if there has been a large correction to spherical aberration during periods of rapid temperature change.

7. The LBC UI doesn't provide any feedback about whether telescope and mirrors are on target. The TCS folks have added this information as a bonus feature on the new IIF GUI. To display this GUI, ssh to lbtmu01 and say "IIFGUI &".

8. With LBC in binocular mode, we normally guide the telescope with the Red side (Master), and we guide the differential motion of the Blue side (Slave) by tilting the primary mirror with active optics commands. Red is chosen as the master since it allows shorter exposures and more guide stars. At present, the Master/Slave selection is hardcoded, and is not controlled by the OB parameters. Blue reverts to being the Master if the Red side is turned off. The hardpoints cause the image to wobble slightly (~1 arcsec) when the mirror position is updated. This wobble makes it undesirable to update the Blue mirror position too often, so we guide the Blue Slave with 16-64 second technical exposures while the Red Master guides the telescope with 4-32 sec exposures. (All these wobble issues will disappear after we upgrade the hardpoints and install fully-binocular PCS later in 2008.) We shortened the minimun Blue guiding exposure from 16 to 8 sec on 7-Feb-2008.

9. Guiding the Blue Slave side by tilting the mirror has a disadvantage that accumulated guiding corrections are not cleared when a new exposure starts. This can cause the co-pointing to drift slightly over time (by a few arcsec). Thus, you may want to ask the operator to "Clear Active Optics" on the Blue/Left side a few times per night. (The operators are trained to Clear Active Optics on both mirrors at the beginning of the night.)

10. An advantage of guiding the Blue Slave side by tilting the mirror is that the SX mirror gets refocussed (according to the collimation lookup table) every time the guiding command is sent. This means that Blue should no longer be limited to 5 minute exposures by drifting focus. Blue U-band exposures of even 10-15 minutes should be possible when guiding in Slave mode.

11. LBC-Blue and SX mirror presently start with about -1000 nm of Z11 from the default configuration. This makes the central hole in the extrafocal pupils hard to identify, and hinders LBCFPIA convergence. A good practice is to ask the telescope operator to apply Z11=+1000 spherical aberration correction to the SX mirror before the first extrafocal image of the night (or whenever Clear Active Optics has been done). (The operators are also trained to do this.)

12. Collimation updates: The primary mirror collimation is updated every time a new preset, dither or focus is commanded. In the rare instance that an observer wishes to take a series of exposures at the same filter without dithering or presetting between them, they should be aware that focus and collimation will not be updated. To force a focus/collimation update, issue a 'stepfocus 0.0' command on the PSF GUI (see issuetrak #363).

13. Rebooting the CMU: If you need to reboot the CMU, be aware that you will also have to run '/home/lbccontrol/lbcstart.sh' to start the LBC daemon before the user interface will start up successfully. This is new - rebooting the CMU used to start the LBC daemon, but for some reason this changed ~winter 2007/8.

14. Beware that the home directories for the new partner accounts are not yet connected to /home/partner, but rather are in /n/nfs/exports/homes/partner which makes it unpleasant to navigate. This is a temporary issue caused by some system reconfiguration to solve another issue. We're sure this will receive prompt attention, because the sysadmin's home directory is in the same condition.

15. The biggest problem with LBC in late-January and early-February has been a problem with the Blue camera detector temperature sensor. When this gets a few bad readings, it makes an error that stops the Blue science camera. A power cycle is required to revive it TurnOff/TurnOn. See Issue #1128 for details.

-- Main.jill - 11 Jan 2007, OlgaKuhn - 07 Jan 2008, JohnHill - 07 Feb 2008
Topic revision: r13 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback